OpenStudy (queenof17):

Solve these proportions. 3/x = x/4 x =

OpenStudy (sunnnystrong):

x=???

alones (alones):

Cross-multiply. \[\frac{ 3 }{ x } = \frac{ x }{4 }\] (3)*(4)=x*x 12=x2 and subtract \(x^{2} \) from both sides 12−\(x^{2} \)=\(x^{2} \)−\(x^{2} \) −\(x^{2} \)+12=0

alones (alones):

subtract 12 from both sides −x^{2}+12−12=0−12 −x^{2}=−12 Than divide both sides by -1 \[\frac{ −x^{2} }{ -1 } = \frac{ -12 }{ -1 }\] \[\ x^{2}=12\]

alones (alones):

Ughh at last take the square root \(x=+2\sqrt{3}\) \(\ x=-2 \sqrt{3}\) Or \[x= \pm2\sqrt{3}\]

OpenStudy (queenof17):

thank you sooo much!!!

alones (alones):

Oh lol my pleasure, i hope that helps. cx